### v1

# Enhanced Athermal Phonon Responsivity in a Kinetic Inductance Detector with Integrated Phonon Collectors

Abstract
Cryogenic phonon detectors are adopted in light dark matter searches and coherent elastic neutrino-nucleus scattering experiments as they can achieve low energy thresholds. The phonon mediated sensing of silicon particle absorbers has already been proved with Kinetic Inductance Detectors (KIDs), acting both as sensors and athermal phonon absorbers at the same time. In this work we present the design and the performance of an improved detector, where the KID acts only as sensor and is coupled to dedicated phonon collectors. The separation between the detector and the collectors increases the variation of the quasi-particle density within the device, thereby enhancing its responsivity. The meander of the KID is composed of a $77\,\mathrm{nm}$ trilayer wire of Aluminum-Titanium-Aluminum, while the phonon collectors are made of a $100\,\mathrm{nm}$ Aluminum layer and act as quasi-particles funnels. Inside the collectors, the absorbed athermal phonons generate quasi-particles which, after diffusion, are trapped in the lower-gap superconducting trilayer. The performance of this setup is compared to that of a standard phonon-mediated KID, showing an increase in responsivity by around a factor of five.
The following article has been submitted to Applied Physics Letters. Copyright 2026 Leonardo Pesce, Alessio Ludovico De Santis, Martino Calvo, Matteo Cappelli, Usasi Chowdhury, Angelo Cruciani, Giorgio Del Castello, Daniele Delicato, Matteo Folcarelli, Matteo del Gallo Roccagiovine, Alessandro Monfardini, Davide Quaranta, and Marco Vignati. This article is distributed under a Creative Commons Attribution (CC BY) License.
Experiments searching for direct light dark matter interaction Abdelhameed et al. (2019); Armengaud et al. (2016); Agnese et al. (2018) or coherent elastic neutrino-nucleus scattering Strauss et al. (2017); Billard et al. (2017); Augier et al. (2023); Agnolet et al. (2017); Akimov et al. (2017) have to address two main technological challenges: the need for high sensitivity to nuclear recoils and the scaling of the instrumented target mass.
The small nuclear recoil energies require detectors with energy thresholds as low as possible. For this reason, cryogenic detectors such as Transition Edge Sensors (TESs) Abdelhameed et al. (2019); Strauss et al. (2017); Alkhatib et al. (2021) or Neutron Transmutation Doped (NTD) Germanium thermistors Armengaud et al. (2016); Augier et al. (2023) are adopted as phonon sensors coupled to target crystals, in order to detect sub-keV nuclear recoil processes. Energy thresholds of the order of $1-10\,\mathrm{eV}$ have already been achieved Anthony-Petersen et al. (2025); Strauss et al. (2017); Alkhatib et al. (2021).
At the same time, the development of mass-scalable instruments is essential for achieving detectors with improved sensitivity to increasingly smaller dark-matter interaction cross sections, assuming a Weakly Interacting Massive Particle (WIMP) model. Experiments with target mass of tens of grams have been built Abdelhameed et al. (2019); Strauss et al. (2017); Armengaud et al. (2019), but scaling-up to the kilogram still remains an experimental challenge.
The BULLKID project developed a particle detector designed for tackling the mass scaling problem Cruciani et al. (2022). It consists of a monolithic array of 60 silicon dices, each one with mass of 0.34 g and volume $5.4\times 5.4\times 5\,\mathrm{mm^{3}}$, carved from a wafer with a diameter of $7.6\,\mathrm{cm}$ and 5 mm thickness. The dices are sensed with 60 KIDs acting as phonon-mediated particle detectors. These are well suited for use in large arrays due to their intrinsic multiplexing capability Day et al. (2003) and therefore can be adopted to design mass-scalable instruments. BULLKID achieved an energy resolution of $27\,\mathrm{eV}$ — which led to set an analysis threshold of around $160~\mathrm{eV}$ Delicato et al. (2024) — that is still far from the current state of art of the other cryogenic detectors. For this reason, it is important to design and develop new sensors aimed at lowering the energy threshold, to the greatest extent possible. In this letter, we investigate a KID design aimed at achieving this goal.
KIDs, which were originally developed for the detection of optical Mazin et al. (2019) and millimeter-wavelength Adam, R. et al. (2018) light, have found application also as phonon-mediated particle detectors Cruciani et al. (2022); Swenson et al. (2010); Cruciani et al. (2012); Golwala et al. (2008). A KID is a superconducting LC resonator Zmuidzinas (2012) in which the total inductance $L$ is the sum of two contributions: the magnetic inductance $L_{g}$, which is related to the geometry of the superconductor, and the kinetic inductance $L_{k}$, which is associated to the Cooper Pairs (CPs) motion inside the metal and depends on their number density. When energy $E$ is released in the substrate onto which the phonon-mediated KID is deposited, athermal phonons are generated and absorbed by the KID itself. If the phonon energy $E_{\mathrm{ph}}$ satisfies $E_{\mathrm{ph}}>2\Delta_{0}$, where $\Delta_{0}\simeq 1.76k_{B}T_{c}$ is the superconducting gap of the metal, $k_{B}$ the Boltzmann constant and $T_{c}$ the critical temperature  Zmuidzinas (2012), CPs in the superconductor are broken and Quasi-Particles (QPs) are generated. This process reduces the CP number density, which in turn increases the kinetic inductance $L_{k}$, leading to a resonant frequency shift, $\Delta f_{r}$, and to a reduction of the resonance depth Zmuidzinas (2012); Mazin (2005); Khalil et al. (2012); Swenson et al. (2013); Vignati et al. (2021).
The change of the CP density is detected by monitoring the phase and magnitude modulation of the complex single pole transmission function $S_{21}(\Delta f_{r})$ of the resonator Delicato et al. (2025):
$S_{21}(\Delta f_{r})=1-\frac{Q}{Q_{c}}\frac{1}{1+2jy}$ (1)
where $Q=(Q_{c}^{-1}+Q_{i}^{-1})^{-1}$ is the total quality factor of the resonator, which combines the coupling quality factor $Q_{c}$ and the internal quality factor $Q_{i}$, that accounts for internal losses due to imperfections and thermal QPs population within the sensor. Finally, the parameter $y=Q\cdot\Delta f_{r}/f_{r}$ is the detuning of the readout frequency relative to the resonance width Zmuidzinas (2012); Mazin (2005); Vignati et al. (2021) Usually, the phase readout $\varphi$ provides a better signal-to-noise ratio Zmuidzinas (2012); therefore, this will be adopted in this work.
The signal responsivity $r$ of a KID in terms of $\varphi$ is Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008); Delicato et al. (2025):
$r=\frac{d\varphi}{dE}=\frac{\eta}{V_{\mathrm{KID}}}\frac{\alpha S_{2}(f_{r},T)Q}{N_{0}\Delta_{0}^{2}}\frac{1}{1+4y^{2}}$ (2)
where $\eta$ is the energy to pair-breaking conversion efficiency, $V_{\mathrm{KID}}$ is the active volume of the sensor, $\alpha=L_{k}/(L_{k}+L_{g})$ is the fraction of kinetic inductance, $S_{2}(f_{r},T)$ is a dimension-less function of order 1 coming from the Mattis-Bardeen theory and $N_{0}$ is the the single-spin density of states. We emphasize that in Eq. 2 the term $\Delta_{0}$ appears quadratically as the product of two identical gap factors, $\Delta_{0}\cdot\Delta_{0}$: the first $\Delta_{0}$ arises from the number of CPs broken by a given energy deposition $E$, which scales as $E/\Delta_{0}$, while the second $\Delta_{0}$ originates from the single–spin volume density, $N_{0}\Delta_{0}$, within the sensor Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008).
If the phonons absorption is not saturated and losses due to supports or interfaces do not dominate, the efficiency $\eta$ scales almost linearly with the phonon collection volume $V_{\mathrm{ph}}$ Martinez et al. (2019); Fyhrie et al. (2016); Santis (2022):
$\eta\propto V_{\mathrm{ph}}$ (3)
In a standard phonon-mediated KID, the volume $V_{\mathrm{ph}}$ coincides with that of the active sensor $V_{\mathrm{KID}}$, since the phonons generated in the substrate are directly absorbed by the inductor. Consequently, from Eq. 2 it follows that the responsivity $r$ of the KID weakly depends on $V_{\mathrm{KID}}$, thus tuning the active sensor volume alone is not expected to lead to a significant improvement in the detector responsivity. In contrast, the separation of these two volumes would allow to design a device with a large $V_{\mathrm{ph}}$ and a small $V_{\mathrm{KID}}$ at the same time, which would enhance the detector responsivity.
We developed and tested a phonon-mediated KID with separate phonon collectors Santis (2022); Mazin et al. (2006); Pesce et al. (2026) that act also as QPs funnels (see Fig. 1). Henceforth, we will refer to this device as the FunKID and to the phonon collectors as the funnels. The resonator is composed of a $77\,\mathrm{nm}$ trilayer meander of Aluminum-Titanium-Aluminum (AlTiAl), with $14\,\mathrm{nm}-33\,\mathrm{nm}-30\,\mathrm{nm}$ thicknesses, respectively Cardani et al. (2018). The inductive part of the KID is composed of seven fingers of around $3600\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}$, evenly spaced by $215\,\mathrm{\mu m}$ (Fig. 1(a)). The funnels are Aluminum structures with $210\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}\times 100\,\mathrm{nm}$ volume intersecting the inductor, evenly spaced by $10\,\mathrm{\mu m}$ (Fig. 1(b)). The total phonon collection volume, which coincides with the total volume of the funnels, is $V_{\mathrm{ph}}\simeq 5.2\,\mathrm{mm}^{2}\times 100\,\mathrm{nm}$, while the active volume of the sensor is $V_{\mathrm{KID}}\simeq 0.55\,\mathrm{mm}^{2}\times 77\,\mathrm{nm}$, giving a volume ratio of $V_{\mathrm{ph}}/V_{\mathrm{KID}}\simeq 12$.
Figure 1: (a) Layout of the FunKID. The funnels in purple intersect the inductive meander indicated with L. The capacity C is adjusted by means of the first two capacitors to tune the resonant frequency. (b) Detail of the funnels. The measures are in $\mathrm{\mu m}$.
Due to proximity effects Cardani et al. (2018); Catalano, A. et al. (2015); Zhao et al. (2017); Brammertz et al. (2001) among the two Aluminum layers ($T_{c}\simeq 1.2\,\mathrm{K}$ Zhao et al. (2017)) and the central Titanium layer ($T_{c}\simeq 0.4\,\mathrm{K}$ Zhao et al. (2017)), the AlTiAl sensor has an intermediate critical temperature $T_{c}$, which is lower than that of the funnels, made of Aluminum. Generally, the AlTiAl critical temperature depends on the boundary resistivities conditions between the layers, and on their thicknesses Zhao et al. (2018). For the same design thicknesses, the CALDER project obtained a value of $T_{c}\simeq 0.81\,\mathrm{K}$ Cardani et al. (2018, 2021), which is expected also for this resonator, provided that the boundary resistance conditions are sufficiently similar.
In this design, schematically reported Fig. 2(a), the athermal phonons produced in the substrate are mostly absorbed by the funnels, where they can break CP generating QPs, which, through diffusion, reach the sensor. Due to the energy gap difference between funnels with gap $\Delta_{\mathrm{Al}}$ and sensor with gap $\Delta_{\mathrm{AlTiAl}}$, QPs with energy $\Delta_{\mathrm{Al}}$ relax down to $\Delta_{\mathrm{AlTiAl}}$ and are trapped within the sensor, preventing their recombination in the funnels Mazin et al. (2006); Riwar and Catelani (2019).
The length of the funnels was chosen to be smaller than the characteristic QP diffusion length in aluminum, $\lambda_{r}=\sqrt{D\tau_{r}}$ Nussbaumer et al. (2000), where $D$ is the QP diffusion coefficient and $\tau_{r}$ the QP recombination time. We adopted a diffusion coefficient of $D=22.5\,\mathrm{cm^{2}/s}$ Hsieh and Levine (1968), a conservative choice compared to higher values reported in the literature Zgirski et al. (2020). The recombination time $\tau_{r}$ depends on both film thickness and temperature, and we adopted $\tau_{r}\simeq 1\,\mathrm{ms}$ de Visser et al. (2011); Barends et al. (2008), yielding a diffusion length of $\lambda_{r}\simeq 1.5\,\mathrm{mm}$, much higher than half the length of the funnels of $100\,\mathrm{\mu m}$. We can also define a characteristic time scale that quantifies the maximum diffusion time of QPs inside the funnels, $\tau_{\mathrm{diff}}\sim l^{2}/D=4\,\mathrm{\mu s}$, where $l=100\,\mathrm{\mu m}$ is half of the funnel length and corresponds to the maximum distance a QP travels before reaching the sensor. This sets the minimum timescale over which the sensor needs to integrate the signal.
We highlight that the funnels have been designed to be as electrically inert as possible, as shown in the simulation performed with the SONNET software Sonnet Software reported in Fig 2(b). In particular, the current of the sensor flows through the inductive meander, while only a small fraction leaks in the funnels. However, we notice that possible limitations of the simulations may arise from the modeling of the overlap region between the funnels and the resonator.
We also notice that in the FunKID the separation between phonon absorbers and sensor modifies the term $\Delta_{0}^{2}$ of Eq. 2. In this design, the CPs breaking occurs in the funnels, with gap $\Delta_{\mathrm{Al}}$, while the QPs population refers to that of the sensor, with gap $\Delta_{\mathrm{AlTiAl}}$. Therefore, the term $\Delta_{0}^{2}$ is replaced by $\Delta_{\mathrm{Al}}\Delta_{\mathrm{AlTiAl}}$.
Figure 2: (a) Detection scheme of the FunKID. Athermal phonons (white dots) are generated by an energy deposition in the substrate. A fraction of these phonons is absorbed by the funnels, which have a higher superconducting gap $\Delta_{\mathrm{Al}}$. There, CPs (red paired dots) are broken, producing QPs (single red dots). Finally, the QPs diffuse into the funnel and are absorbed by the KID, which features a lower gap $\Delta_{\mathrm{AlTiAl}}$, where they recombine. (b) Current density simulated with SONNET. Most of the current flows into the KID, while a small leakage is observed. The length reported is measured in $\mathrm{\mu m}$.
According to the simulation performed with SONNET, the FunKID is designed to have a small coupling quality factor $Q_{c}=20\,\mathrm{k}$ and resonant frequency $f_{r}=1009.6\,\mathrm{MHz}$, corresponding to an expected ring time $\tau_{\mathrm{ring}}=Q/\pi f_{r}=5\,\mathrm{\mu s}$ Cruciani et al. (2022). This is the integration time of the sensor and thus should be higher than the diffusion time $\tau_{\mathrm{diff}}$, in order to maximize the signal in the sensor. In our case, these time constants are comparable, negatively impacting on the amplitude of the signal. However, the choice of operating with small quality factor lies on possible limitations of the internal quality factor $Q_{i}$ Zmuidzinas (2012); Mazin et al. (2006), which not only would limit the total quality factor $Q$, but also would broaden the width and decrease the depth of the resonance, as follows from Eq. 1.
We estimated the gain in responsivity of the FunKID relative to an identical standard KID without funnels, $r_{\mathrm{FunKID}}/r_{\mathrm{KID}}$, using the Eq. 2:
$\frac{r_{\mathrm{FunKID}}}{r_{\mathrm{KID}}}\simeq\frac{V_{\mathrm{ph}}}{V_{\mathrm{KID}}}\frac{\alpha_{\mathrm{FunKID}}}{\alpha_{\mathrm{KID}}}\frac{Q_{\mathrm{FunKID}}}{Q_{\mathrm{KID}}}\frac{\Delta_{\mathrm{Al}}}{\Delta_{\mathrm{AlTiAl}}}$ (4)
where we assumed that $N_{0}$ and $S_{2}(f_{r},T)$ were approximately equal among the two devices and that their response was linear, thus $1+4y^{2}\simeq 1$. Moreover, we considered only the geometric dependence of $\eta$ for both the KIDs according to Eq. 3 and that the phonon collection volume of the standard KID coincides with its active volume $V_{\mathrm{KID}}$. In order to evaluate the best gain in responsivity, we assumed same quality factors $Q$ and kinetic inductance fractions $\alpha$ for both the devices and the critical temperatures of Aluminum and AlTiAl of $1.2\,\mathrm{K}$ and $0.81\,\mathrm{K}$, respectively. Under these conditions, the expected gain amounts to about 8.
The test device was fabricated at the PTA facility clean room in Grenoble Plateforme technologique amont using a $500\,\mathrm{\mu m}$ thick silicon tile. On the same substrate, a standard phonon-mediated KID was produced along with the FunKID (Fig. 3). The KID sensors were etched from a trilayer of Aluminum ($30\,\mathrm{nm}$), Titanium ($14\,\mathrm{nm}$) and Aluminum ($13\,\mathrm{nm}$) following a standard lithographic technique. After the fabrication of the KIDs, photo-resist was used to coat the surface of the wafer, and gaps in the shape of the funnels were patterned on it. The funnels were then realized by selectively depositing $100\,\mathrm{nm}$ of Aluminum in the patterned gaps, following a lift-off process. The two resonators are identical except for the first two capacitors length: in the FunKID these were designed $105\,\mathrm{\mu m}$ longer to lower its resonant frequency by approximately $20\,\mathrm{MHz}$ with respect to the KID. Finally, we highlight that the feed-line is fabricated from the same AlTiAl trilayer used for the meanders of the two KIDs.
Figure 3: The two pixels on the silicon tile suspended by Teflon supports. On the left is the FunKID, while on the right the standard phonon-mediated KID.
The tile was mounted into a copper holder by means of Teflon supports and wire bonded to $50\,\mathrm{\Omega}$ launchers. The device was operated in a dilution $^{3}\mathrm{He}/^{4}\mathrm{He}$ cryostat, at a base temperature $T_{0}\simeq 35\,\mathrm{mK}$, much lower than the AlTiAl critical temperature. The copper holder was closed and covered by aluminum foil to mitigate environmental radiation. Finally, a mu-metal shield was placed outside the cryostat to reduce magnetic-field effects Flanigan et al. (2016).
To calibrate the detectors and characterize their pulses, three optical fibers were placed facing the side of the silicon tile opposite to the KIDs: two positioned to face the center of each resonator, and one located midway between them. The two lateral fibers were used to identify the two pixels, while the central one was used to calibrate the KIDs from the same position for comparison. The fibers were pulsed with $400\,\mathrm{nm}$ photons bursts generated at room temperature by controlled a LED lamp Del Castello (2024).
The readout tones were generated by the NIXA electronic system Bourrion et al. (2011, 2013), a superheterodyne radio-frequency detection system capable to transmit and receive tones at the KIDs resonant frequencies. The transmitted waves undergo an attenuation chain of $55\,\mathrm{dB}$ before the chip, are transmitted past the resonators, then are amplified by $30\,\mathrm{dB}$ by a Low Noise Amplifier (model: CITLF3 Cosmic Microwave Technology, Inc. (2025)) operated at $4\,\mathrm{K}$ and finally received back by NIXA.
The resonances of the two KIDs are shown in Fig. 4. We fitted both the amplitude of the transmission function $|S_{21}|$ and its real and imaginary components in order to extract the resonant frequencies of the two KIDs as well as the quality factors $Q_{c}$, $Q_{i}$, and $Q$. We also evaluated the resonator ring times $\tau_{\mathrm{ring}}=Q/\pi f_{r}$ of the two KIDs. The results are reported in Tab. 1.
Figure 4: Transmission function amplitude $|S_{21}|$ around the resonance frequencies $f_{r}$ of the resonators.
$f_{r}\,\mathrm{[MHz]}$ $Q_{c}\,\mathrm{[k]}$ $Q_{i}\,\mathrm{[k]}$ $Q\,\mathrm{[k]}$ $\tau_{\mathrm{ring}}\,\mathrm{[\mu s]}$
FunKID 968.6 19 20 10 3
KID 1029.0 20 410 19 6
Table 1: Results of the resonant frequency $f_{r}$, quality factors $Q$, $Q_{i}$ and $Q_{c}$ and ring times $\tau_{\mathrm{ring}}$ of the two resonators. The uncertainty of the resonant frequency is negligible, while the relative uncertainty on the internal quality factor $Q_{i}$ was estimated to around $10\,\%$.
A shift of around $40\,\mathrm{MHz}$ of the measured FunKID resonant frequency with respect to the simulation was observed, indicating that the overlap regions between the funnels and the resonator might have not been correctly modeled in the simulations. Conversely, the obtained KID resonant frequency is in agreement with simulation, assuming a kinetic inductance $L_{k}=1.4\,\mathrm{pH/sq}$, which is the same of Ref. Cardani et al. (2018). The coupling quality factors $Q_{c}\sim 20\,\mathrm{k}$ are in line with the values simulated with SONNET, for both resonators. In addition, the internal quality factor of the FunKID is comparable to the coupling quality factor, which currently limits this design. Conversely, the internal quality of the KID is higher than the FunKID and in line with the values obtained in Ref. Cardani et al. (2018), indicating that the reduced $Q_{i}$ of the FunKID is set by the funnels and not by fabrication processes of the trilayer AlTiAl film.
For each resonator, we evaluated our best estimate of kinetic inductance fraction $\alpha$, by combining the measured resonant frequencies $f_{r}$ with the resonant frequencies $f_{0}$ obtained from SONNET simulations performed with a perfect conductor (i.e., with $L_{k}=0$), through the formula:
$\alpha=1-\left(\frac{f_{r}}{f_{0}}\right)^{2}$ (5)
The simulations were calibrated against two identical resonators but in Aluminum. We obtained $\alpha=18.9\,\%$ and $\alpha=15.6\,\%$ for the FunKID and the KID, respectively, much larger than those obtained in Aluminum, $\alpha=7.0\,\%$ and $\alpha=5.0\,\%$ using the same design, consistently with an increase of kinetic inductance $L_{k}$ Cardani et al. (2018).
We also measured the feed-line critical temperature $T_{c}=(0.83\pm 0.08)\,\mathrm{K}$, corresponding to a superconducting gap $\Delta_{0}=(125.9\pm 12.6)\,\mathrm{\mu eV}$, and we assumed that this coincided with the critical temperature of the normal KID. Such value is in reasonable agreement with that found by the CALDER project Cardani et al. (2018) of $0.81\,\mathrm{K}$, using a similar film. Nevertheless such estimation is reliable as long as it is applied to the normal KID. The presence of funnels is expected to modify the FunKID critical temperature, but in the current setup we do not have a direct way to measure it.
We also measured $\delta f_{r}/f_{r}$, the relative shift of the resonant frequency with respect to the base temperature resonant frequency, $f_{r}$. Using a model based on the BCS theory Gao et al. (2008), it is possible to write:
$\frac{\delta f_{r}}{f_{r}}=-\frac{\alpha}{2}S_{2}(f_{r},T)\sqrt{\frac{{2\pi k_{B}T}}{\Delta(T)}}e^{-\Delta(T)/k_{B}T}$ (6)
where $\Delta(T)$ is the AlTiAl superconducting order parameter Gao et al. (2008). For this purpose, we measured the resonant frequency of each resonator at different temperatures, while keeping the transmission bias powers fixed at $87\,\mathrm{dBm}$ and $71\,\mathrm{dBm}$ for the FunKID and the KID, respectively. These powers are lower than the operative calibration powers in order to isolate the effect of thermally generated excess QPs in the resonators, while keeping the micro-wave induced contribution as low as possible de Visser et al. (2012). However, potential limitations may arise when applying such a model to inhomogeneous and multilayer superconductors. In order to test the applicability of such model to our devices, we compared the data with the BCS model calculated with our best predictions of $\alpha$ and $\Delta_{0}$ (Fig. 5).
Figure 5: Relative shift of the resonant frequency with respect to the base temperature resonant frequency, $\delta f_{r}/f_{r}$, for the FunKID in blue and KID in purple. The dashed curves are obtained with the BCS model calculated with our best predictions of $\alpha$ and $\Delta_{0}$.
In both cases, data deviate from the BCS model. The FunKID shows significant discrepancies at intermediate temperatures, indicating that the presence of the funnels and the trilayer KID structure prevents the BCS model from fully describing the device. For the normal KID, the overall behavior of data is similar to the BCS model, although deviations at high temperatures are still observed. Therefore more complete models than the BCS-based one would be required for a more accurate interpretation Zhao et al. (2017), but these are beyond the scope of this work.
To calibrate the detectors, the FunKID and the KID were operated at bias transmission calibration powers $P_{\mathrm{{cal}}}$ of $68\,\mathrm{dBm}$ and $62\,\mathrm{dBm}$, respectively, at which they exhibited good performance in terms of signal to noise ratio Cardani et al. (2018). The choice to operate the FunKID with a lower calibration power was motivated by the observation that on the FunKID non-linear effects became relevant at smaller power than the KID Swenson et al. (2013). Optical pulses of increasing total energy from $\sim 9\,\mathrm{keV}$ up to $\sim 90\,\mathrm{keV}$ were fired onto the back of the substrate by means of the central fiber to characterize the response of the two KIDs Del Castello (2024). Finally, the signals were processed offline with a matched filter Gatti and Manfredi (1986) in order to further improve the signal to noise ratio.
Two pulses generated by an energy release of around $90\,\mathrm{keV}$ into the substrate have been superimposed in Fig. 6. It is possible to observe that the FunKID exhibits a pulse response about 5.5 times higher than the KID. Our best prediction for the gain in responsivity under identical electrical conditions was around 8. According to Eq. 4 the dominant degradation of the FunKID response might be associated with its smaller quality factor $Q$, which is a factor 1.9 lower than that of the KID. However, our best prediction of $\alpha$ of the FunKID is about 1.2 higher than that of the KID. By correcting the best gain prediction with these two effects, we obtain a gain of around 5.1, similar to measurements.
The decay times $\tau_{d}$ of the FunKID and KID are about $0.15\,\mathrm{ms}$ and $0.45\,\mathrm{ms}$, respectively. The longer decay time of the KID is in line with the higher internal quality factor Cruciani et al. (2016). The rise times $\tau_{\mathrm{rise}}$ amount to around $12\,\mathrm{\mu s}$ and $20\,\mathrm{\mu s}$, respectively. These are significantly higher than the respective ring times (see Tab 1). In addition, the rise time of the FunKID is higher than the QPs diffusion time scale $\tau_{\mathrm{diff}}\simeq 4\,\mathrm{\mu s}$ previously defined. It seems reasonable that in both cases the rise times of the pulses are dominated by the phonon absorption time $\tau_{\mathrm{ph}}$ Swenson et al. (2010); Martinez et al. (2019), which can be faster in the case of the FunKID due to its higher phonon collection efficiency.
Figure 6: Two pulses generated by the same energy release. A higher response is observed for the FunKID resonator compared to the normal KID. Also, it is possible to notice the higher $\tau_{d}$ of the KID pulse, which takes almost $2\,\mathrm{ms}$ to come back to the baseline level.
The results of the energy calibrations Cardani et al. (2018); Delicato et al. (2025); Del Castello (2024) of the two KIDs, the phase responsivity $r$, the noise RMS after applying the matched filter, $\sigma_{N}$, and the corresponding baseline energy resolution of the device, $\sigma_{E}$, have been reported in Tab. 2. The phase responsivity is about 5.3 higher at the calibration power for the FunKID, which is in line with the measured responsivity gain previously discussed. Moreover, an improvement of around a factor of 2 for the energy resolution in $\mathrm{eV}$ is measured, despite the worse noise RMS $\sigma_{N}$ of the FunKID. The improvement of the FunKID responsivity was therefore proven, even with a total quality factor smaller by a factor of about 2.
FunKID KID
$P_{\mathrm{{cal}}}\,\mathrm{[dBm]}$ -68 -62
$r\,\mathrm{[mrad/keV]}$ $5.85\pm 0.14$ $1.11\pm 0.04$
$\sigma_{N}\,\mathrm{[mrad]}$ $0.386\pm 0.009$ $0.173\pm 0.005$
$\sigma_{E}\,\mathrm{[eV]}$ $65.9\pm 2.3$ $156\pm 6$
Table 2: Results of the energy calibrations performed on the two KIDs. For each device, we report the calibration power $P_{\mathrm{cal}}$, the phase responsivity $r$, the noise RMS after the matched filter $\sigma_{N}$ and energy resolution $\sigma_{E}$.
While the tested device served as a proof of concept and the improvement in detector responsivity was demonstrated, an enhancement in energy resolution is still required, as the current performance remains worse than previously achieved results Delicato et al. (2024). Achieving this improvement will demand further optimization, such as an additional increase of the total quality factors $Q$, which in turn requires higher internal quality factors $Q_{i}$.
Acknowledgements.
This work was supported by the INFN, Sapienza University of Rome and co-funded by the European Union (ERC, DANAE, 101087663). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We thank A. Girardi and M. Iannone of the INFN Sezione di Roma for technical support. We acknowledge the support of the PTA platform for the fabrication of the device.
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.