### v2

# Enhanced Athermal Phonon Responsivity in a Kinetic Inductance Detector with Integrated Phonon Collectors

Abstract
Cryogenic phonon detectors are adopted in light dark matter searches and coherent elastic neutrino-nucleus scattering experiments as they can achieve low energy thresholds. The phonon mediated sensing of silicon particle absorbers has already been proved with Kinetic Inductance Detectors (KIDs), acting both as sensors and athermal phonon absorbers at the same time. In this work we present the design and the performance of an improved detector, where the KID acts only as sensor and is coupled to dedicated phonon collectors. The separation between the detector and the collectors increases the variation of the quasi-particle density within the device, thereby enhancing its responsivity. The meander of the KID is composed of a $77\,\mathrm{nm}$ trilayer wire of Aluminum-Titanium-Aluminum, while the phonon collectors are made of a $100\,\mathrm{nm}$ Aluminum layer and act as quasi-particles funnels. Inside the collectors, the absorbed athermal phonons generate quasi-particles which, after diffusion, are trapped in the lower-gap superconducting trilayer. The performance of this setup is compared to that of a standard phonon-mediated KID, showing an increase in responsivity by around a factor of five.
The following article has been submitted to Applied Physics Letters. Copyright 2026 Leonardo Pesce, Alessio Ludovico De Santis, Martino Calvo, Matteo Cappelli, Usasi Chowdhury, Angelo Cruciani, Giorgio Del Castello, Daniele Delicato, Matteo Folcarelli, Matteo del Gallo Roccagiovine, Alessandro Monfardini, Davide Quaranta, and Marco Vignati. This article is distributed under a Creative Commons Attribution (CC BY) License.
Experiments searching for direct light dark matter interaction Abdelhameed et al. (2019); Armengaud et al. (2016); Agnese et al. (2018) or coherent elastic neutrino-nucleus scattering Strauss et al. (2017); Billard et al. (2017); Augier et al. (2023); Agnolet et al. (2017); Akimov et al. (2017) address two main technological challenges: the need for high sensitivity to nuclear recoils and the scaling of the instrumented target mass.
The small nuclear recoil energies require detectors with energy thresholds as low as possible. For this reason, cryogenic detectors Abdelhameed et al. (2019); Armengaud et al. (2016); Strauss et al. (2017); Augier et al. (2023); Alkhatib et al. (2021) are adopted as phonon sensors coupled to target crystals, while energy thresholds of the order of 1-10 eV have already been achieved Anthony-Petersen et al. (2025); Strauss et al. (2017); Alkhatib et al. (2021). At the same time, the development of mass-scalable instruments is essential for achieving detectors with improved sensitivity to increasingly smaller dark-matter interaction cross sections, assuming a Weakly Interacting Massive Particle (WIMP) model Roszkowski et al. (2018). Experiments with target mass of tens of grams have been built Abdelhameed et al. (2019); Strauss et al. (2017); Armengaud et al. (2019), but scaling-up to the kilogram still remains an experimental challenge.
The BULLKID project developed a particle detector designed for tackling the mass scaling problem Cruciani et al. (2022). It consists of a monolithic array of 60 silicon dices, each one with mass of 0.34 g and volume $5.4\times 5.4\times 5\,\mathrm{mm^{3}}$, carved from a wafer with a diameter of 7.6 cm and 5 mm thickness. The dices are sensed with 60 KIDs acting as phonon-mediated particle detectors. The best energy resolution achieved amounts to 27 eV — which led to set an analysis threshold of around 160 eV Delicato et al. (2024) — which is still far from the current state of art of the other cryogenic detectors. For this reason, it is important to develop new sensors aimed at lowering the energy threshold, to the greatest extent possible. In this letter, we investigate a KID design aimed at achieving this goal.
KIDs, which were originally developed for the detection of optical Mazin et al. (2019) and millimeter-wavelength Adam, R. et al. (2018) light, have found application also as phonon-mediated particle detectors Cruciani et al. (2022); Swenson et al. (2010); Cruciani et al. (2012); Golwala et al. (2008). A KID is a superconducting LC resonator Zmuidzinas (2012) in which the total inductance $L$ is the sum of the magnetic inductance $L_{g}$ and the kinetic inductance $L_{k}$. When energy $E$ is released in the substrate of the KID, athermal phonons are generated and absorbed by the KID, where they can break Cooper Pairs (CPs) releasing Quasi-Particles (QPs). This process reduces the CP number density, which in turn increases the kinetic inductance $L_{k}$, leading to a resonant frequency shift, $\Delta f_{r}$, and to a reduction of the resonance depth Zmuidzinas (2012); Mazin (2005); Khalil et al. (2012); Swenson et al. (2013); Vignati et al. (2021).
The change of the CP density is detected by monitoring the phase and magnitude modulation of the complex single pole transmission function $S_{21}(y)$ of the resonator Delicato et al. (2025):
$S_{21}(y)=1-\frac{Q}{Q_{c}}\frac{1}{1+2jy}$ (1)
where $Q=(Q_{c}^{-1}+Q_{i}^{-1})^{-1}$ is the total quality factor of the resonator, which combines the coupling quality factor $Q_{c}$ and the internal quality factor $Q_{i}$ Zmuidzinas (2012). Finally, the parameter $y=Q\cdot\Delta f_{r}/f_{r}$ is the detuning of the readout frequency relative to the resonance width Zmuidzinas (2012); Mazin (2005); Vignati et al. (2021). Usually, the phase readout $\varphi$ provides a better signal-to-noise ratio Zmuidzinas (2012); therefore, this will be adopted in this work.
The signal phase responsivity $r$ of a KID is Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008); Delicato et al. (2025):
$r=\frac{\eta}{V_{\mathrm{KID}}}\frac{\alpha S_{2}(f_{r},T)Q}{N_{0}\Delta_{0}^{2}}\frac{1}{1+4y^{2}}$ (2)
where $\eta$ is the energy to pair-breaking conversion efficiency, $V_{\mathrm{KID}}$ is the active inductor volume, $\alpha=L_{k}/(L_{k}+L_{g})$ is the kinetic inductance fraction, and $S_{2}(f_{r},T)$ is a dimensionless factor of order one from Mattis-Bardeen theory. Finally, $N_{0}$ is the single-spin density of states and $\Delta_{0}\simeq 1.76k_{B}T_{c}$ is the superconducting gap, with $k_{B}$ the Boltzmann constant and $T_{c}$ the critical temperature. We emphasize that in Eq. 2 the term $\Delta_{0}$ appears quadratically as the product of two identical gap factors: one term arises from the number of CPs broken by a given energy deposition $E$, scaling as $E/\Delta_{0}$, while the second originates from the single–spin volume density, $N_{0}\Delta_{0}$, within the sensor Zmuidzinas (2012); Mazin (2005); Moore et al. (2012); Gao et al. (2008).
If phonons absorption is not saturated and losses due to supports or interfaces do not dominate, the efficiency $\eta$ scales almost linearly with the phonon collection volume $V_{\mathrm{ph}}$ Martinez et al. (2019); Fyhrie et al. (2016); Santis (2022):
$\eta\propto V_{\mathrm{ph}}$ (3)
In a standard phonon-mediated KID, where $V_{\mathrm{ph}}$ coincides with the active volume $V_{\mathrm{KID}}$, the responsivity weakly depends on the inductor volume, as follows from Eq. 2. Conversely, separating these volumes allows to design devices with large $V_{\mathrm{ph}}$ and small $V_{\mathrm{KID}}$, which would enhance the detector responsivity.
We developed and tested a phonon-mediated KID with separate phonon collectors Santis (2022); Mazin et al. (2006); Pesce et al. (2026) that act also as QPs funnels (see Fig. 1). We refer to this device as the FunKID and to the phonon collectors as the funnels. The resonator is composed of a 77 nm trilayer meander of Aluminum-Titanium-Aluminum (AlTiAl), with 14 nm-33 nm-30 nm thicknesses, respectively Cardani et al. (2018). The inductive part of the KID is composed of seven fingers of around $3600\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}$, evenly spaced by $215\,\mathrm{\mu m}$ (Fig. 1(a)). The funnels are Aluminum structures with $210\,\mathrm{\mu m}\times 10\,\mathrm{\mu m}\times 100\,\mathrm{nm}$ volume intersecting the inductor, evenly spaced by $10\,\mathrm{\mu m}$ (Fig. 1(b)). The total phonon collection volume, which coincides with the total volume of the funnels, is $V_{\mathrm{ph}}\simeq 5.2\,\mathrm{mm}^{2}\times 100\,\mathrm{nm}$, while the active volume of the sensor is $V_{\mathrm{KID}}\simeq 0.55\,\mathrm{mm}^{2}\times 77\,\mathrm{nm}$, giving a volume ratio of $V_{\mathrm{ph}}/V_{\mathrm{KID}}\simeq 12$.
Figure 1: (a) Layout of the FunKID. The funnels in purple intersect the inductive meander indicated with L, while the capacitor is indicated with C. (b) Detail of the funnels. The measures are in $\mathrm{\mu m}$.
Due to proximity effects Cardani et al. (2018); Catalano, A. et al. (2015); Zhao et al. (2017); Brammertz et al. (2001) among the two Aluminum layers ($T_{c}\simeq 1.2\,\mathrm{K}$ Zhao et al. (2017)) and the central Titanium layer ($T_{c}\simeq 0.4\,\mathrm{K}$ Zhao et al. (2017)), the AlTiAl sensor has an intermediate critical temperature $T_{c}$, which is lower than that of the funnels, made of Aluminum. For the same design thicknesses, the CALDER project measured the critical temperature of an AlTiAl trilayer as $T_{c}\simeq 0.81\,\mathrm{K}$ Cardani et al. (2018, 2021); a similar value is expected for this resonator, assuming comparable boundary effects Zhao et al. (2018).
In the FunKID design, schematically reported Fig. 2(a), the phonons produced in the substrate are mostly absorbed by the funnels, where they can break CP generating QPs, which, through diffusion, reach the sensor. Due to the energy gap difference between funnels with gap $\Delta_{\mathrm{Al}}$ and sensor with gap $\Delta_{\mathrm{AlTiAl}}$, QPs with energy $\Delta_{\mathrm{Al}}$ relax down to $\Delta_{\mathrm{AlTiAl}}$ and are trapped within the sensor, preventing their recombination in the funnels Mazin et al. (2006); Riwar and Catelani (2019).
The length of the funnels was chosen to be smaller than the characteristic QP diffusion length in aluminum, $\lambda_{r}=\sqrt{D\tau_{r}}$ Nussbaumer et al. (2000), where $D$ is the QP diffusion coefficient and $\tau_{r}$ the QP recombination time. We adopted $D=22.5\,\mathrm{cm^{2}/s}$ Hsieh and Levine (1968) and $\tau_{r}\simeq 1\,\mathrm{ms}$ de Visser et al. (2011); Barends et al. (2008), yielding to a diffusion length of $\lambda_{r}\simeq 1.5\,\mathrm{mm}$, much higher than half the length of the funnels of $100\,\mathrm{\mu m}$. We can also define a characteristic time scale quantifying the maximum diffusion time of QPs inside the funnels before reaching the sensor, $\tau_{\mathrm{diff}}\sim l^{2}/D=4\,\mathrm{\mu s}$, where $l=100\,\mathrm{\mu m}$ is half of the funnel length. This sets the minimum timescale over which the sensor needs to integrate the signal.
We highlight that the funnels have been designed to be as electrically inert as possible, as shown in the simulation performed with the SONNET software Sonnet Software reported in Fig 2(b). In particular, the current of the sensor flows through the inductive meander, while only a small fraction leaks in the funnels. However, we notice that possible limitations of the simulations may arise from the modeling of the overlap region between the funnels and the resonator.
We also notice that in the FunKID the separation between phonon absorbers and sensor modifies the term $\Delta_{0}^{2}$ of Eq. 2. In this design, the CPs breaking occurs in the funnels, with gap $\Delta_{\mathrm{Al}}$, while the QPs population refers to that of the sensor, with gap $\Delta_{\mathrm{AlTiAl}}$. Therefore, the term $\Delta_{0}^{2}$ is replaced by $\Delta_{\mathrm{Al}}\Delta_{\mathrm{AlTiAl}}$.
Figure 2: (a) Detection scheme of the FunKID. The white dots represent the athermal phonons generated by an energy deposition in the substrate, the red paired dots are the CPs and the single red dots the QPs diffusing from the funnel to the KID. More details in the text. (b) Current density simulated with SONNET. The length reported is measured in $\mathrm{\mu m}$.
According to the simulation performed with SONNET, the FunKID is designed to have a small coupling quality factor $Q_{c}=20\,\mathrm{k}$ and resonant frequency $f_{r}=1009.6\,\mathrm{MHz}$, corresponding to an expected ring time $\tau_{\mathrm{ring}}=Q/\pi f_{r}=5\,\mathrm{\mu s}$ Cruciani et al. (2022). This is the integration time of the sensor and thus should be higher than the diffusion time $\tau_{\mathrm{diff}}$, in order to maximize the signal in the sensor. In our case, these time constants are comparable, negatively impacting on the amplitude of the signal. However, the choice of operating with small quality factor lies on possible limitations of the internal quality factor $Q_{i}$ Zmuidzinas (2012); Mazin et al. (2006), which not only would limit the total quality factor $Q$, but also would broaden the width and decrease the depth of the resonance, as follows from Eq. 1.
The responsivity gain of the FunKID relative to an identical standard KID without funnels (with phonon collection volume $V_{\mathrm{KID}}$), $r_{\mathrm{FunKID}}/r_{\mathrm{KID}}$, is found using Eq. 2:
$\frac{r_{\mathrm{FunKID}}}{r_{\mathrm{KID}}}\simeq\frac{V_{\mathrm{ph}}}{V_{\mathrm{KID}}}\frac{\alpha_{\mathrm{FunKID}}}{\alpha_{\mathrm{KID}}}\frac{Q_{\mathrm{FunKID}}}{Q_{\mathrm{KID}}}\frac{\Delta_{\mathrm{AlTiAl}}}{\Delta_{\mathrm{Al}}}$ (4)
where we assumed that $N_{0}$ and $S_{2}(f_{r},T)$ were approximately equal between the two devices and that their response was linear, thus $1+4y^{2}\simeq 1$. Furthermore, we considered only the geometric dependence of $\eta$ for both KIDs, according to Eq. 3. Under identical electrical conditions, indeed equal quality factors $Q$ and kinetic inductance fractions $\alpha$ for both devices we expect a responsivity gain of approximately 8.
Figure 3: The two pixels on the silicon tile suspended by Teflon supports. On the left is the FunKID, while on the right the standard phonon-mediated KID.
The test device was fabricated at the PTA facility clean room in Grenoble Plateforme technologique amont using a $500\,\mathrm{\mu m}$ thick silicon tile. On the same substrate, an identical standard phonon-mediated KID, with a simulated resonant frequency $20\,\mathrm{MHz}$ higher than the FunKID, was also produced (Fig. 3). The KID sensors and the feed-line were etched from the 77 nm AlTiAl trilayer following a standard lithographic technique. After their fabrication, photo-resist was used to coat the surface of the wafer, and gaps in the shape of the funnels were patterned on it. The funnels were then realized by selectively depositing $100\,\mathrm{nm}$ of Aluminum in the patterned gaps, following a lift-off process.
The tile was mounted into a copper holder by means of Teflon supports and wire bonded to $50\,\mathrm{\Omega}$ launchers. The device was operated in a dilution $^{3}\mathrm{He}/^{4}\mathrm{He}$ cryostat, at a base temperature $T_{0}\simeq 35\,\mathrm{mK}$. Finally, a mu-metal shield was placed outside the cryostat to reduce magnetic-field effects Flanigan et al. (2016).
To calibrate the detectors, three optical fibers were placed facing the substrate backside: two were aligned with the resonators for their identification Pesce et al. (2026) and one was located at the center to calibrate the KIDs from the same position for comparison. The fibers were pulsed with $400\,\mathrm{nm}$ photons bursts generated at room temperature by a controlled LED lamp Del Castello (2024).
The readout tones were generated by the NIXA superheterodyne radio-frequency detection system Bourrion et al. (2011, 2013) transmitting and receiving tones at the KIDs resonant frequencies. The transmitted waves undergo an attenuation chain of $55\,\mathrm{dB}$ before the chip, are transmitted past the resonators, then are amplified by 30 dB by a Low Noise Amplifier Cosmic Microwave Technology, Inc. (2025) operated at 4 K and finally received back by NIXA.
The resonances of the two KIDs are shown in Fig. 4. We fitted both the amplitude of the transmission function $|S_{21}|$ and its real and imaginary components in order to extract the resonant frequencies and the quality factors $Q_{c}$, $Q_{i}$, and $Q$ of the two KIDs. We also evaluated their resonator ring times $\tau_{\mathrm{ring}}=Q/\pi f_{r}$. The results are reported in Tab. 1.
Figure 4: Transmission function amplitude $|S_{21}|$ around the resonance frequencies $f_{r}$ of the resonators.
$f_{r}\,\mathrm{[MHz]}$ $Q_{c}\,\mathrm{[k]}$ $Q_{i}\,\mathrm{[k]}$ $Q\,\mathrm{[k]}$ $\tau_{\mathrm{ring}}\,\mathrm{[\mu s]}$
FunKID 968.6 19 20 10 3
KID 1029.0 20 410 19 6
Table 1: Results of the resonant frequency $f_{r}$, quality factors $Q$, $Q_{i}$ and $Q_{c}$ and ring times $\tau_{\mathrm{ring}}$ of the two resonators.
A shift of around $40\,\mathrm{MHz}$ of the measured FunKID resonant frequency with respect to the simulation was observed, indicating that the overlap regions between the funnels and the resonator might have not been correctly modeled in the simulations. Conversely, the obtained KID resonant frequency is in agreement with simulation, assuming a kinetic inductance $L_{k}=1.4\,\mathrm{pH/sq}$, which is the same of Ref. Cardani et al. (2018). The coupling quality factors $Q_{c}\sim 20\,\mathrm{k}$ are in line with the values simulated with SONNET, for both resonators. The internal quality factor of the FunKID is comparable to the coupling quality factor, which currently limits this design. Conversely, the internal quality of the KID is higher than the FunKID and in line with the values obtained in Ref. Cardani et al. (2018), indicating that the reduced $Q_{i}$ of the FunKID is set by the funnels and not by fabrication processes of the trilayer AlTiAl film.
For each resonator, we evaluated an estimate of kinetic inductance fraction $\alpha$, by combining the measured resonant frequencies $f_{r}$ with the resonant frequencies $f_{0}$ obtained from SONNET simulations performed with a perfect conductor ($L_{k}=0$Cardani et al. (2018):
$\alpha=1-\left(\frac{f_{r}}{f_{0}}\right)^{2}$ (5)
The simulations were calibrated against two identical resonators but in Aluminum. We obtained $\alpha=18.9\,\%$ and $\alpha=15.6\,\%$ for the FunKID and the KID, respectively, much larger than those obtained in Aluminum, $\alpha=7.0\,\%$ and $\alpha=5.0\,\%$ using the same design.
We measured a feed-line $T_{c}=(0.83\pm 0.08)\,\mathrm{K}$, corresponding to $\Delta_{0}=(125.9\pm 12.6)\,\mathrm{\mu eV}$, in reasonable agreement with the $0.81\,\mathrm{K}$ reported in Cardani et al. (2018). We assumed that this value applies to both the resonators, although funnels likely modify the FunKID critical temperature, which cannot be measured directly in the presented setup.
We measured the resonant frequency shift in temperature relative to the base temperature resonant frequency, $\delta f_{r}/f_{r}$, with bias powers fixed at $87\,\mathrm{dBm}$ (FunKID) and $71\,\mathrm{dBm}$ (KID), chosen to minimize the microwave-induced QPs excess de Visser et al. (2012). We compared the data to a model based on the BCS theory Gao et al. (2008), calculated with the predicted $\alpha$ and $\Delta_{0}$ previously discussed (Fig. 5). However, since the BCS theory might be limited in case of inhomogeneous and multilayer superconductors, the comparison is qualitative and aims to test the applicability of the model.
Figure 5: Relative shift of the resonant frequency with respect to the base temperature resonant frequency, $\delta f_{r}/f_{r}$, for the FunKID in blue and KID in purple. The dashed curves are obtained with the BCS model calculated with our best predictions of $\alpha$ and $\Delta_{0}$.
In both cases, data deviate from the BCS model. The FunKID shows significant discrepancies at intermediate temperatures, indicating that the presence of the funnels and the trilayer KID structure prevents the BCS model from fully describing the device. For the normal KID, the overall behavior of data is similar to the BCS model, although deviations at high temperatures are still observed. Therefore more complete models than the BCS-based one would be required for a more accurate interpretation Zhao et al. (2017), but these are beyond the scope of this work.
To calibrate the detectors, the FunKID and the KID were operated at bias transmission calibration powers $P_{\mathrm{{cal}}}$ of -68 dBm and -62 dBm, respectively, at which they exhibited good performance in terms of signal to noise ratio Cardani et al. (2018). Optical pulses of increasing total energy from around 9 keV up to 90 keV were fired onto the back of the substrate by means of the central fiber Del Castello (2024). Finally, signals were processed offline with a matched filter Gatti and Manfredi (1986) in order to further improve the signal to noise ratio.
Two pulses generated by an energy release of around 90 keV into the substrate have been superimposed in Fig. 6. The FunKID exhibits a pulse response about 5.5 times larger than the standard KID. This value is lower than the prediction of 8 derived assuming identical electrical conditions. However, by correcting this estimate using Eq. 4 with the measured ratios of quality factors and kinetic inductance fractions (approximately 0.53 and 1.21, respectively), we obtain a gain of around 5.1, similar to measurements. The decay times $\tau_{d}$ of the FunKID and KID are about 0.15 ms and 0.45  ms, respectively. The smaller decay time of the FunKID is in line with the smaller internal quality factor Cruciani et al. (2016). The rise times $\tau_{\mathrm{rise}}$ amount to around 12 $\mathrm{\mu s}$ and 20 $\mathrm{\mu s}$, respectively. These are significantly higher than the respective ring times (see Tab 1). In addition, the rise time of the FunKID is higher than the QPs diffusion time scale $\tau_{\mathrm{diff}}\simeq 4\,\mathrm{\mu s}$ previously defined. It seems reasonable that in both cases the rise times of the pulses are dominated by the phonon absorption time $\tau_{\mathrm{ph}}$ Swenson et al. (2010); Martinez et al. (2019), which can be faster in the case of the FunKID due to its higher phonon collection efficiency.
Figure 6: Two pulses generated by the same energy release. A higher response is observed for the FunKID resonator compared to the normal KID.
The results of the energy calibrations Cardani et al. (2018); Delicato et al. (2025); Del Castello (2024) have been reported in Tab. 2. The responsivity is about 5.3 higher at the calibration power for the FunKID, which is in line with the measured responsivity gain previously discussed. Moreover, an improvement of around a factor of 2 for the energy resolution in eV is measured, despite the worse noise RMS $\sigma_{N}$ of the FunKID. The improvement of the FunKID responsivity was therefore proven, even with a smaller total quality factor.
FunKID KID
$P_{\mathrm{{cal}}}\,\mathrm{[dBm]}$ -68 -62
$r\,\mathrm{[mrad/keV]}$ $5.85\pm 0.14$ $1.11\pm 0.04$
$\sigma_{N}\,\mathrm{[mrad]}$ $0.386\pm 0.009$ $0.173\pm 0.005$
$\sigma_{E}\,\mathrm{[eV]}$ $65.9\pm 2.3$ $156\pm 6$
Table 2: Results of the energy calibrations performed on the two KIDs. For each device, we report the calibration power $P_{\mathrm{cal}}$, the phase responsivity $r$, the noise RMS after the matched filter $\sigma_{N}$ and energy resolution $\sigma_{E}$.
While the tested device served as a proof of concept and the improvement in detector responsivity was demonstrated, an enhancement in energy resolution is still required, as the current performance remains worse than previously achieved results Delicato et al. (2024). Achieving this improvement demands further optimization, such as an additional increase of the total quality factors $Q$, which in turn requires higher internal quality factors $Q_{i}$.
Acknowledgements.
This work was supported by the INFN, Sapienza University of Rome and co-funded by the European Union (ERC, DANAE, 101087663). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We thank A. Girardi and M. Iannone of the INFN Sezione di Roma for technical support. We acknowledge the support of the PTA platform for the fabrication of the device.
Author declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.